# Towards Scalable Newborn Screening: Automated General Movement Assessment in Uncontrolled Settings

### v3

@@ 3.2 Results
| Coord. | |||||||
| #1 | Angles | ||||||
| Both | |||||||
| Coord. | |||||||
| #2 | Angles | ||||||
| Both | |||||||
| Coord. | |||||||
| #3 | Angles | ||||||
| Both | |||||||
| Coord. | 0.6460 | 0.5699 | |||||
| Mean | Angles | 0.7717 | 0.5892 | 0.5440 | |||
| Both | 0.7484 | ||||||
| Coord. | |||||||
| Automatic | Angles | ||||||
| Both | |||||||
Table 2 shows the results of the different classification models, for each labeller individually as well as averaged and labeled with AggPose, measured in AUROC. See Appendix B for a comprehensive overview of AUPRC and accuracy in each task. They demonstrate a variety of outcomes across age groups and feature sets, with no clear pattern indicating that one method is superior. However, the use of angle features, either isolated or combined, proves useful. Additionally, the early GM group shows significantly better overall results, showing a big difference in the difficulty of the task across age groups. Overall, the Label & Track approach performs comparably to the automatically labeled AggPose method. However, in a significant number of data groups, it demonstrates improved results, highlighting the positive impact of keypoint tracking. By incorporating temporal information from the time series, this approach enhances classification performance, whereas AggPose labels each keypoint in isolated frames, missing potential contextual cues. One of the key takeaways from this study is that despite the relatively small and heterogeneous dataset, our approach has shown success in automatically classifying GM in newborns, with average AUROCs of up to 0.8283 in certain scenarios, opening promising lines of research for GMA “in the wild”.
4 Conclusion and Future Outlook
In conclusion, our study demonstrates that we can successfully predict GMA with a promising AUROC of up to 0.8283, with performance generally better for the early GM group. This highlights the potential of using automated methods to assess infant motor behaviour. Despite the challenges posed by diverse data sources and recording conditions in the data, our results show that automated GMA is feasible in real-world clinical settings. This is a significant step forward in the early detection of neurodevelopmental disorders. This study lays the groundwork for further advancements, emphasizing the need for larger datasets and more refined models and feature extraction approaches to capture the data complexities. Ultimately, the goal is to enhance early prediction of neurodevelopmental diseases, offering a valuable tool for clinicians, and ultimately improving patient care.
One of the key weaknesses of this study is the relatively small dataset, which limits the robustness and generalizability of our results, as indicated by relatively large standard deviations across folds. Moreover, the heterogeneity of the data makes the problem more challenging, while broadening its impact in the clinics. Expanding the dataset would allow us to explore more sophisticated methods, such as Multiple Instance Learning (MIL), which could introduce interpretability by making predictions on a per-episode basis. Similarly, using Vision Transformers to directly process the videos could help better capture spatial and temporal patterns, further improving classification performance. Future work within the current dataset could focus on developing further features from the extracted keypoints and exploring other model classes to leverage the temporal structure in the task.
Acknowledgments
DC is funded from the Strategic Focal Area “Personalized Health and Related Technologies (PHRT)” grant #2021-911 of the ETH Domain (Swiss Federal Institutes of Technology), and SL and KC from the Swiss State Secretariat for Education, Research and Innovation (SERI), contract #MB22.00047. The authors would like to thank Ričards Marcinkevičs and Heike Leutheuser for their insights and discussions on the development of this manuscript.
Appendix A Extended Dataset Description
In this section, we present additional details about our dataset to the best of our ability. Some details are not available to us, such as weight at video reconding time for early GM group infants. Table 3 shows several demographic characteristics, as well as details about the videos themselves for both the early and late GM groups. Figure 6 provides histograms for the distributions of both video resolution and FPS for the early and late GM groups.
| Characteristic | Early GM Group | Late GM Group |
| Age at video recording (Days) (mean stddev) | 170.4 24.8 | |
| Post-menstrual age at video recording (Days) (mean stddev) | 231.5 6.5 | N/A |
| Corrected Age at video recording (Days) (mean stddev) | N/A | 92.5 17.4 |
| Weight at video recording [g] (mean, stddev) | N/A | 4148.2 830.7 |
| Video Length [s] (mean stddev) | 409 144 | 224 93 |
| FPS [min, max] | [30, 120] | [30, 120] |
| Extremely Preterm | 16 | 15 |
| Very Preterm | 20 | 19 |
| Moderate to Late Preterm | 3 | 3 |
| Not Preterm | 0 | 0 |
Appendix B Further results
Table 4 and Table 5 show the results of the presented classification models, for each labeller individually as well as averaged, grouped by age and feature extraction method used, measured in AUPRC and accuracy, respectively. The best performing method per task is marked in bold and overall both metrics resemble the behaviour of the AUROCs reported in the main text in Table 2.
| Labeller | Features | Early GM Group | Late GM Group | ||||
| AUPRC | AUPRC | ||||||
@@ Appendix B Further results
| Coord. | |||||||
| 1 | Angles | ||||||
| Both | |||||||
| Coord. | |||||||
| 2 | Angles | ||||||
| Both | |||||||
| Coord. | |||||||
| 3 | Angles | ||||||
| Both | |||||||
| Coord. | 0.7429 | 0.8366 | 0.4981 | ||||
| Mean | Angles | 0.8415 | 0.4498 | 0.4236 | |||
| Both | |||||||
| Coord. | |||||||
| Automatic | Angles | ||||||
| Both | |||||||
| Labeller | Features | Early GM Group | Late GM Group | ||||
| Accuracy | Accuracy | ||||||
| RF | LSTM | CNN | RF | LSTM | CNN | ||
| Coord. | |||||||
| 1 | Angles | ||||||
| Both | |||||||
| Coord. | |||||||
| 2 | Angles | ||||||
| Both | |||||||
| Coord. | |||||||
| 3 | Angles | ||||||
| Both | |||||||
| Coord. | |||||||
| Mean | Angles | 0.6926 | 0.6688 | 0.6438 | |||
| Both | 0.6272 | 0.6742 | 0.6225 | ||||
| Coord. | |||||||
| Automatic | Angles | ||||||
| Both | |||||||

### v4

@@ Appendix B Further Results
| AUROC | |||||||
| Coord. | 0.1005* | 0.1226* | 0.2578 | ||||
| Angles | 0.2641 | 0.2546 | 0.1511 | -0.1146* | -0.1625 | ||
| Both | 0.2122 | 0.1091* | 0.2383 | -0.1737* | |||
| AUPRC | |||||||
| Coord. | 0.0890* | 0.0954 | 0.1865 | -0.1412 | |||
| Angles | 0.1623 | 0.1500 | -0.1529 | -0.2629 | |||
| Both | 0.1587 | 0.0857 | 0.1591 | -0.2397 | |||
| Accuracy | |||||||
| Coord. | 0.0540 | 0.1980 | 0.0939 | ||||
| Angles | 0.1912* | 0.2000* | 0.0504 | -0.0693* | -0.0876 | ||
| Both | 0.1271 | 0.0987* | 0.1863 | ||||
| Metric | Input Features | Early GM Group | Late GM Group | ||||
@@
| AUROC | |||||||
| Coord. | 0.0213* | 0.0016* | 0.0000 | ||||
| Angles | 0.0000 | 0.0000 | 0.0075 | 0.0192* | 0.0051 | ||
| Both | 0.0000 | 0.0069* | 0.0000 | 0.0004* | |||
| AUPRC | |||||||
| Coord. | 0.0070* | 0.0040 | 0.0000 | 0.0030 | |||
| Angles | 0.0000 | 0.0000 | 0.0040 | 0.0000 | |||
| Both | 0.0000 | 0.0107 | 0.0000 | 0.0000 | |||
| Accuracy | |||||||
| Coord. | 0.0405 | 0.0000 | 0.0048 | ||||
| Angles | 0.0000* | 0.0000* | 0.0325 | 0.0248* | 0.0071 | ||
| Both | 0.0008 | 0.0072* | 0.0000 | ||||